# Believing without Seeing: Quality Scores for Contextualizing Vision-Language Model Explanations

### v2

@@ Abstract
Keyu He1   Tejas Srinivasan2   Brihi Joshi2 Xiang Ren2Jesse Thomason2Swabha Swayamdipta2 1Carnegie Mellon University 2University of Southern California keyuhe@cmu.edu
@@ 2.1 Visual Fidelity
We measure the visual fidelity of an explanation $E$ with respect to the image $I$ by decomposing the explanation into a set of facts about the image and individually verifying each fact (Algorithm 1). Conceptually, this is an oracle scoring function that would check each atomic claim against ground-truth visual Q&A; we approximate it with an LLM-as-judge, using a verifier $m_{\texttt{Verif}}$ (GPT-4o by default), and we audit this approximation with expert annotations and verifier swaps at Appendix˜J. Concretely, we prompt a question generation LLM $m_{\texttt{QGen}}$ to generate a set of verification questions $\mathcal{Q}^{V}$ that confirm visual details that have been mentioned in the VLM-generated explanation. The model has to generate questions such that answering “yes” confirms the presence of the detail in the image (see Table˜7 for the full LLM prompt). The number of questions generated by $m_{\texttt{QGen}}$ depends on the explanation and may vary across instances.
@@ 3 Are Visual Fidelity and Contrastiveness Indicative of VLM Correctness?
We evaluate our explanation qualities on three visual reasoning tasks: A-OKVQA Schwenk et al. (2022)—a multiple-choice VQA benchmark that requires reasoning over images using external knowledge and commonsense; VizWiz Gurari et al. (2018)—an open-ended VQA task consisting of questions asked by blind and low-vision users about images they captured on their mobile phones; and MMMU-Pro Yue et al. (2025)—a massive multi-discipline multimodal understanding and reasoning benchmark. We sample 500 questions from the validation set of each task. Section˜K.2 contains details about the tasks’ pre-processing and evaluation. For VizWiz, which lacks a specified set of possible answers, we generate hard negatives using an LLM to enable Contrastiveness evaluation. These negatives are used for post-hoc evaluation and are not provided to the model during answer generation. [^2] We discuss the specific challenges and results of this open-set evaluation in Appendix F and validate the strategy via ablation in Appendix G.
@@
Table 1: Discriminability (Disc., higher is better) for different explanation quality scoring functions across three models and three visual reasoning tasks. Significance is evaluated using an unpaired t-test: ∗∗∗ $p<0.001$, ∗∗ $p<0.01$, $p<0.05$. Averages are not evaluated for significance. Notably, VizWiz–GPT-4o exhibits significantly lower and predominantly negative discriminability scores across all metrics; see Appendix˜I for a discussion of the underlying causes.
@@ Results.
We observe that across all datasets and models, our proposed quality scoring functions almost always achieve higher Discriminability scores and lower Expected Calibration Errors than existing scoring functions across all models. Most differences in Discriminability on these datasets are statistically significant ($p<0.001$), confirming that our metrics effectively distinguish accurate from inaccurate predictions. We also find that our explanation-based metrics significantly outperform standard logit-based uncertainty baselines on complex reasoning tasks (see Appendix˜H). On VizWiz, while LLaVA-v1.5-7B shows strong discriminability, GPT-4o and Qwen2.5-VL-7B exhibit lower or negative scores on all scoring functions; as discussed in Appendix˜I, this is primarily due to valid answers being penalized by the dataset’s strict ground truth. Crucially, combining the Visual Fidelity and Contrastiveness quality scores results in the best overall Discriminability and Calibration, highlighting their complementary utility.
@@
To test whether VF’s behavior depends on the specific verifier model, we replace the default GPT-4o with two smaller open verifiers (Gemma-3n-E4B and Qwen2.5-VL-7B). Section˜J.3 shows that all three yield comparable discriminability and calibration, indicating VF remains stable across verifier architectures.
@@ 4.1 User Study Setup
We conducted a between-subjects study on a balanced subset of 100 questions each from A-OKVQA and VizWiz (total 300 annotations per condition). Participants were recruited via Prolific and paid \$2 base + performance bonuses (see Appendix˜K for full protocol and demographics).
@@ 4.2 RQ1: Does providing explanation quality scores improve user reliance?
We first evaluate whether showing users our proposed quality scores alongside VLM explanations helps users more accurately assess VLM correctness. We compare communicating: 1) only the Visual Fidelity score, 2) only the Contrastiveness score, 3) both scores, and 4) a product of the two scores. Figure 9 contains examples of how these quality scores are communicated to users. We additionally include a control setting where only an explanation is shown without a quality score, and a random setting where users are shown a quality score randomly sampled between 0 and $1$.
@@ 5 Conclusions and Future Directions
One limitation of our approach lies in the open-ended Contrastiveness evaluation. Its score quality is inherently partially bounded by the semantic quality of the generated hard negatives. While we validate that our generator produces challenging distractors in Appendix˜I, the metric remains sensitive to their plausibility; if distractors are too easily distinguished, the metric may overestimate confidence, whereas if they are semantically ambiguous, it may penalize valid reasoning. Thus, our metric’s resolution is coupled with the generator’s capability. Second, while we demonstrate effectiveness on A-OKVQA, its multiple-choice format may not strictly be representative of real-world, open-ended visual queries. We attempt to bridge this by evaluating on open-ended VizWiz examples and the diverse MMMU-Pro benchmark, but future work should explore even broader task distributions. Finally, we only evaluate on English-language datasets and conduct user studies with fluent English speakers. Future work should explore multilingual generalization.
@@
This work was conducted with careful consideration of ethical implications. Human studies were conducted with informed consent through Prolific, with participants compensated at rates exceeding US minimum wage (\$2 base + performance bonuses, see Section˜K.3.2). Participants were native English speakers from the US with higher education credentials (Table˜18). We acknowledge this may limit generalizability across demographic groups.
@@
The VizWiz dataset used contains questions from blind and low-vision users but is fully anonymized. In our human studies, the annotators were anonymized. We filtered NSFW content (Section˜K.2) and did not collect any additional personal information. All model outputs were manually reviewed for offensive content before presentation to study participants.
@@ Acknowledgments
This research is supported in part by the National Science Foundation (NSF) under grant IIS2403437, the Simons Foundation, and the Allen Institute for AI. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of NSF. B. Joshi was supported by the Apple Scholars in AI/ML PhD Fellowship. This work was partially done when S. Swayamdipta was a visitor at the Simons Institute for the Theory of Computing.
@@ Appendix H Uncertainty Quantification Baselines
To situate our explanation quality metrics within broader uncertainty quantification literature, we compare them against a standard logit-based baseline. Specifically, for the LLaVA-v1.5-7B model (where we have access to output probabilities), we compute the softmax probability of the first token of predicted answer as a measure of model confidence.
@@
Table 12: Comparison of our explanation quality metrics against a Logit-based Baseline (softmax probability of the predicted answer’s first token) on LLaVA-v1.5-7B. Significance is evaluated using an unpaired t-test: ∗∗∗ $p<0.001$, ∗∗ $p<0.01$, $p<0.05$.
Table˜12 presents the comparison on two contrasting datasets: A-OKVQA (simpler visual reasoning) and MMMU-Pro (complex, multi-discipline reasoning).
On A-OKVQA, the model’s intrinsic “intuition” (logits) is highly effective, achieving strong Discriminability ($0.252$) and excellent ECE ($0.046$). This suggests that for simpler tasks, raw confidence is a sufficient reliability indicator.
On MMMU-Pro, however, this signal breaks down. Logits yield poor discriminability ($0.055$, not significant) and high calibration error ($0.356$). In contrast, our explanation quality metrics—particularly the product of Visual Fidelity and Contrastiveness ($S_{\mathrm{Prod}}$)—remain robust, achieving statistically significant discriminability ($0.090^{**}$) and much better calibration ($0.144$).
@@
From Table˜13, we show our evaluation framework on four representative VQA instances. For each example, we provide the original question (and answer choices where applicable), image, ground truth answer, the model’s selected answer and rationale, and the corresponding scores for Simulatability, Informativeness, Commonsense Plausibility, Visual Fidelity, and Contrastiveness. Note that for VizWiz, the ground-truth answer is defined as the majority vote among the 10 user annotations from the original dataset, and since VizWiz provides no answer choices, we do not compute a Contrastiveness score.
@@
Appendix I VizWiz Dataset Limitations and Negative Discriminability
@@
Upon qualitative inspection, we found this phenomenon is largely driven by the “Ground Truth Problem” inherent in the VizWiz dataset. The ground truth answers are often narrow or specific, causing valid model predictions to be labeled as incorrect. Table˜14 shows three representative examples:
@@
Appendix J Reliability and Validity of the Trust Cues
J.1 Validation on Agreement with Human Judgements
To validate the reliability of our automated Visual Fidelity and Contrastiveness qualities, we conducted an internal sanity check on a held‐out sample. We randomly selected 100 examples from our evaluation sets and had a single expert annotator manually assess each instance. The instruction provided to the expert annotator is shown at Table˜17. We binarized our scores at a threshold of $0.5$ (scores $\geq 0.5$ map to $1$, else 0) and compared these labels to the human judgments using Cohen’s $\kappa$:
$\kappa_{\mathrm{VF}}=\,\,0.51\quad\text{and}\quad\kappa_{\mathrm{Contr.}}=\,\,0.44,$
indicating moderate agreement.
J.2 Accuracy of LLM-generated verification questions and answers
@@
J.3 Verifier robustness for Visual Fidelity
@@
As shown in Table˜15 and Table˜16, across A-OKVQA and VizWiz, all three verifiers yield positive discriminability on average (GPT-4o: 0.109; Gemma-3n-E4B: 0.061; Qwen2.5-VL-7B: 0.154), and ECE remains comparably low (0.168/0.197/0.196 respectively). Thus, VF continues to separate correct from incorrect predictions and stays well-calibrated even with using smaller open models as verifiers.
@@
Appendix K User Studies
K.1 Annotator filtering
@@
To ensure high-quality annotations, we recruited annotators on Prolific using the criteria listed in Table˜18. All participants identities were fully anonymized. All annotators provided informed consent through Prolific’s platform, which clearly described the task purpose, data use, and compensation prior to participation.
This study protocol was reviewed and approved by our institution’s Institutional Review Board (the name of the institution will be included in the camera-ready version).
@@
K.2 Dataset Splits and User Study Design
@@
These subsets were balanced to contain 50 correct and 50 incorrect predictions by LLaVA-v1.5-7B. To ensure the subsets remained representative of the larger distribution, we generated 50 random candidate subsets and selected the one with the lowest Expected Calibration Error (ECE) relative to the full 500-instance set. Specifically, for A-OKVQA, we minimized the average ECE across Visual Fidelity and Contrastiveness; for VizWiz, we minimized the ECE for Visual Fidelity (as Contrastiveness scores were unavailable during the subset creation phase). As shown in Table˜20 and Table˜20, this selection maintains a distribution of quality scores consistent with the full set from Section˜3.
@@
In the study, each question was annotated by three unique participants, yielding 300 annotations per dataset per condition. Each participant completed a batch of 10 questions, balanced to include five instances where the model was correct and five where it was incorrect; participants were not informed of this distribution. Users were randomly assigned to a single study condition (control or a treatment variant) and did not participate in multiple conditions. In total, we collect 300 human annotations per dataset and condition (3 users × 100 questions), capturing user reliance behavior under varying explanation and quality signal conditions.
@@
K.4 User Studies Settings & Examples
@@
[…]
@@
In the main user study, we present 14 settings in total. Table˜21 summarizes all 14 different settings and Figure˜9 shows examples of all 8 different categories of messages presented to users. Table˜22 contains the main user-study results of user behavior patterns.
K.5 Supplementary Human Studies
In addition to the study described in Section˜4.2 and Section˜K.4, we ran a supplementary user study to evaluate how explanations and qualities affect users’ trust in model predictions. Each of the 10 questions was presented in three successive stages; after each stage, participants indicated whether they believed the model’s answer was correct or were unsure:
@@
Participants were paid a \$2 base fee and could earn up to \$1 in performance‐based bonuses, which were awarded only during Stage 3 (Explanation + Quality; see Section˜K.3.2).
@@
Table˜23 shows that as users progress from seeing only the model’s answer to viewing explanations and then explanations with quality scores, their unsure rate steadily decreases while user accuracy correspondingly increases, with the largest gains observed when descriptive qualities are provided alongside explanations. These findings support our findings that richer, more interpretable quality signals can meaningfully improve users’ trust calibration.
Appendix L Computational Resources Spent & Total Cost
The complete breakdown of all monetary expenditures across the study is given in Table˜24. In brief, the computational expenses include six hours of A100 GPU time for running the vision‐language models, plus OpenAI API calls for generating predicted answers, explanations, and computing qualities. Human evaluation costs cover two Prolific studies—a one‐stage study as shown in Section˜K.4 and a three‐stage follow‐up from Section˜K.5 (including up to \$1.00 in per‐participation bonuses; see Section˜K.3.2).

### v3

@@ Abstract
Keyu He1,2   Tejas Srinivasan1   Brihi Joshi1 Xiang Ren1Jesse Thomason1Swabha Swayamdipta1 1University of Southern California 2Carnegie Mellon University keyuhe@cmu.edu
@@ 2.1 Visual Fidelity
We measure the visual fidelity of an explanation $E$ with respect to the image $I$ by decomposing the explanation into a set of facts about the image and individually verifying each fact (Algorithm 1). Conceptually, this is an oracle scoring function that would check each atomic claim against ground-truth visual Q&A; we approximate it with an LLM-as-judge, using a verifier $m_{\texttt{Verif}}$ (GPT-4o by default), and we audit this approximation with expert annotations and verifier swaps at Appendix˜K. Concretely, we prompt a question generation LLM $m_{\texttt{QGen}}$ to generate a set of verification questions $\mathcal{Q}^{V}$ that confirm visual details that have been mentioned in the VLM-generated explanation. The model has to generate questions such that answering “yes” confirms the presence of the detail in the image (see Table˜7 for the full LLM prompt). The number of questions generated by $m_{\texttt{QGen}}$ depends on the explanation and may vary across instances.
@@ 3 Are Visual Fidelity and Contrastiveness Indicative of VLM Correctness?
We evaluate our explanation qualities on three visual reasoning tasks: A-OKVQA Schwenk et al. (2022)—a multiple-choice VQA benchmark that requires reasoning over images using external knowledge and commonsense; VizWiz Gurari et al. (2018)—an open-ended VQA task consisting of questions asked by blind and low-vision users about images they captured on their mobile phones; and MMMU-Pro Yue et al. (2025)—a massive multi-discipline multimodal understanding and reasoning benchmark. We sample 500 questions from the validation set of each task. Section˜L.2 contains details about the tasks’ pre-processing and evaluation. For VizWiz, which lacks a specified set of possible answers, we generate hard negatives using an LLM to enable Contrastiveness evaluation. These negatives are used for post-hoc evaluation and are not provided to the model during answer generation. [^2] We discuss the specific challenges and results of this open-set evaluation in Appendix F and validate the strategy via ablation in Appendix G.
@@
Table 1: Discriminability (Disc., higher is better) for different explanation quality scoring functions across three models and three visual reasoning tasks. Significance is evaluated using an unpaired t-test: ∗∗∗ $p<0.001$, ∗∗ $p<0.01$, $p<0.05$. Averages are not evaluated for significance. Notably, VizWiz–GPT-4o exhibits significantly lower and predominantly negative discriminability scores across all metrics; see Appendix˜J for a discussion of the underlying causes.
@@ Results.
We observe that across all datasets and models, our proposed quality scoring functions almost always achieve higher Discriminability scores and lower Expected Calibration Errors than existing scoring functions across all models. Most differences in Discriminability on these datasets are statistically significant ($p<0.001$), confirming that our metrics effectively distinguish accurate from inaccurate predictions. We also find that our explanation-based metrics significantly outperform standard logit-based uncertainty baselines on complex reasoning tasks (see Appendix˜H). On VizWiz, while LLaVA-v1.5-7B shows strong discriminability, GPT-4o and Qwen2.5-VL-7B exhibit lower or negative scores on all scoring functions; as discussed in Appendix˜J, this is primarily due to valid answers being penalized by the dataset’s strict ground truth. Crucially, combining the Visual Fidelity and Contrastiveness quality scores results in the best overall Discriminability and Calibration, highlighting their complementary utility.
@@
To test whether VF’s behavior depends on the specific verifier model, we replace the default GPT-4o with two smaller open verifiers (Gemma-3n-E4B and Qwen2.5-VL-7B). Section˜K.3 shows that all three yield comparable discriminability and calibration, indicating VF remains stable across verifier architectures.
@@ 4.1 User Study Setup
We conducted a between-subjects study on a balanced subset of 100 questions each from A-OKVQA and VizWiz, with 30 participants and 300 annotations per dataset per condition. Participants were recruited via Prolific and paid \$2 base + performance bonuses (see Appendix˜L for full protocol and demographics).
@@ 4.2 RQ1: Does providing explanation quality scores improve user reliance?
We first evaluate whether showing users our proposed quality scores alongside VLM explanations helps users more accurately assess VLM correctness. We compare communicating: 1) only the Visual Fidelity score, 2) only the Contrastiveness score, 3) both scores, and 4) a product of the two scores. Figure 9 contains examples of how these quality scores are communicated to users. We additionally include a control setting where only an explanation is shown without a quality score, and a random setting where users are shown a quality score drawn uniformly at random from $[0,1]$, presented with the same UI as the “simple confidence” variant in Figure˜9(d).
@@ 5 Conclusions and Future Directions
One limitation of our approach lies in the open-ended Contrastiveness evaluation. Its score quality is inherently partially bounded by the semantic quality of the generated hard negatives. While we validate that our generator produces challenging distractors in Appendix˜J, the metric remains sensitive to their plausibility; if distractors are too easily distinguished, the metric may overestimate confidence, whereas if they are semantically ambiguous, it may penalize valid reasoning. Thus, our metric’s resolution is coupled with the generator’s capability. Second, while we demonstrate effectiveness on A-OKVQA, its multiple-choice format may not strictly be representative of real-world, open-ended visual queries. We attempt to bridge this by evaluating on open-ended VizWiz examples and the diverse MMMU-Pro benchmark, but future work should explore even broader task distributions. Finally, we only evaluate on English-language datasets and conduct user studies with fluent English speakers. Future work should explore multilingual generalization.
@@
This work was conducted with careful consideration of ethical implications. Human studies were conducted with informed consent through Prolific, with participants compensated at rates exceeding US minimum wage (\$2 base + performance bonuses, see Section˜L.3.2). Participants were native English speakers from the US with higher education credentials (Table˜19). We acknowledge this may limit generalizability across demographic groups.
@@
The VizWiz dataset used contains questions from blind and low-vision users but is fully anonymized. In our human studies, the annotators were anonymized. We filtered NSFW content (Section˜L.2) and did not collect any additional personal information. All model outputs were manually reviewed for offensive content before presentation to study participants.
@@ Acknowledgments
This research is supported in part by the National Science Foundation (NSF) under grant IIS2403437, the Simons Foundation, the Allen Institute for AI, DARPA award HR00112490376, and a Capital One CREDIF research grant. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of NSF. B. Joshi was supported by the Apple Scholars in AI/ML PhD Fellowship. This work was partially done when S. Swayamdipta was a visitor at the Simons Institute for the Theory of Computing.
@@ Appendix H Uncertainty Quantification Baselines
To situate our explanation quality metrics within broader uncertainty quantification literature, we compare them against two standard uncertainty baselines. Specifically, for the LLaVA-v1.5-7B model (where we have access to output probabilities), we evaluate:
  1. 1.
    Model Confidence (Logits): the softmax probability of the first token of the predicted answer.
  2. 2.
    $P(\text{True})$: prompting the model “Is the proposed answer correct?” and computing a softmax over the logits of affirmative versus negative tokens.
Method Prompt
$P(\text{True})$ Question: {question}
Choices: {choices}
Proposed answer: {predicted_answer}
Is the proposed answer correct? Respond with only "True" or "False".
Table 12: Prompt used for the $P(\text{True})$ uncertainty quantification baseline. We extract the logits of affirmative tokens and negative tokens at the first generation step, taking the maximum within each group, and normalize via softmax to obtain $P(\text{True})$.
The exact prompts are shown in Table˜12.
@@
Table 13: Comparison of our explanation quality metrics against uncertainty quantification baselines on LLaVA-v1.5-7B. We consider these baseline metrics: logit-based confidence (softmax probability of the predicted answer’s first token) and $P(\text{True})$ (probability assigned to “True” when asked if the prediction is correct). Significance is evaluated using an unpaired t-test: ∗∗∗ $p<0.001$, ∗∗ $p<0.01$, $p<0.05$.
Table˜13 presents the comparison on two contrasting datasets: A-OKVQA (simpler visual reasoning) and MMMU-Pro (complex, multi-discipline reasoning).
On A-OKVQA, the model’s intrinsic “intuition” (logits) is highly effective, achieving strong Discriminability ($0.252$) and excellent ECE ($0.046$). P(True) ($0.032$) exhibits substantially weaker discriminability and poor calibration (ECE >0.25>0.25). This suggests that for simpler tasks, raw confidence is a sufficient reliability indicator.
On MMMU-Pro, both uncertainty baselines break down. Logits yield poor discriminability ($0.055$, not significant) and high calibration error ($0.356$); P(True) produces especially poor calibration (ECE of $0.772$). In contrast, our explanation quality metrics—particularly the product of Visual Fidelity and Contrastiveness ($S_{\mathrm{Prod}}$)—remain robust, achieving statistically significant discriminability ($0.090^{**}$) and much better calibration ($0.144$).
@@
Appendix I Qualitative Examples
Table˜14 shows our evaluation framework on four representative VQA instances. For each example, we provide the original question (and answer choices where applicable), image, ground truth answer, the model’s selected answer and rationale, and the corresponding scores for Simulatability, Informativeness, Commonsense Plausibility, Visual Fidelity, and Contrastiveness. Note that for VizWiz, the ground-truth answer is defined as the majority vote among the 10 user annotations from the original dataset.
@@ Appendix I Qualitative Examples
Appendix J VizWiz Dataset Limitations and Negative Discriminability
@@
Upon qualitative inspection, we found this phenomenon is largely driven by the “Ground Truth Problem” inherent in the VizWiz dataset. The ground truth answers are often narrow or specific, causing valid model predictions to be labeled as incorrect. Table˜15 shows three representative examples:
@@
Appendix K Reliability and Validity of the Trust Cues
K.1 Validation on Agreement with Human Judgements
To validate the reliability of our automated Visual Fidelity and Contrastiveness qualities, we conducted an internal sanity check on a held‐out sample. We randomly selected 100 examples from our evaluation sets and had two expert annotators independently assess each instance. The first annotator is one of the authors of this paper, who has direct experience in VLM evaluation; the second is an independent expert who also has experience with VLM evaluation but is not familiar with this paper or the dataset. The instruction provided to the expert annotators is shown at Table˜18.
We binarized our scores at a threshold of $0.5$ (scores $\geq 0.5$ map to $1$, else 0) and computed Cohen’s $\kappa$ both between the two annotators and between each annotator and the automated scores. Inter-annotator agreement between the two experts is
$\kappa_{\mathrm{VF}}=0.68\quad\text{and}\quad\kappa_{\mathrm{Contr.}}=0.46,$
confirming that the task is well-defined. Agreement with the automated scores, reported as (Annotator 1 / Annotator 2), is
$\kappa_{\mathrm{VF}}=0.51\,/\,0.49\quad\text{and}\quad\kappa_{\mathrm{Contr.}}=0.44\,/\,0.39,$
indicating moderate and consistent agreement across annotators and between annotators and automated qualities.
K.2 Accuracy of LLM-generated verification questions and answers
@@
K.3 Verifier robustness for Visual Fidelity
@@
As shown in Table˜16 and Table˜17, across A-OKVQA and VizWiz, all three verifiers yield positive discriminability on average (GPT-4o: 0.109; Gemma-3n-E4B: 0.061; Qwen2.5-VL-7B: 0.154), and ECE remains comparably low (0.168/0.197/0.196 respectively). Thus, VF continues to separate correct from incorrect predictions and stays well-calibrated even with using smaller open models as verifiers.
@@
Appendix L User Studies
L.1 Annotator filtering
@@
To ensure high-quality annotations, we recruited annotators on Prolific using the criteria listed in Table˜19. All participants identities were fully anonymized. All annotators provided informed consent through Prolific’s platform, which clearly described the task purpose, data use, and compensation prior to participation.
This study protocol was reviewed and approved by the University of Southern California’s Institutional Review Board.
@@
L.2 Dataset Splits and User Study Design
@@
These subsets were balanced to contain 50 correct and 50 incorrect predictions by LLaVA-v1.5-7B. To ensure the subsets remained representative of the larger distribution, we generated 50 random candidate subsets and selected the one with the lowest Expected Calibration Error (ECE) relative to the full 500-instance set. Specifically, for A-OKVQA, we minimized the average ECE across Visual Fidelity and Contrastiveness; for VizWiz, we minimized the ECE for Visual Fidelity (as Contrastiveness scores were unavailable during the subset creation phase). As shown in Table˜21 and Table˜21, this selection maintains a distribution of quality scores consistent with the full set from Section˜3.
@@
In the study, each question was annotated by three unique participants, yielding 30 participants and 300 annotations per dataset per condition. Each participant completed a batch of 10 questions, balanced to include five instances where the model was correct and five where it was incorrect; participants were not informed of this distribution. Users were randomly assigned to a single study condition (control or a treatment variant) and did not participate in multiple conditions. In total, we collect 300 human annotations per dataset and condition (3 users × 100 questions), capturing user reliance behavior under varying explanation and quality signal conditions.
@@
L.4 User Studies Settings & Examples
@@
(d) Simple confidence score – Numeric quality (used for Prod(VF, Contr.) and Random Score)
@@
In the main user study, we present 14 settings in total. Table˜22 summarizes all 14 different settings and Figure˜9 shows examples of all 8 different categories of messages presented to users. Table˜23 contains the main user-study results of user behavior patterns.
L.5 Supplementary Human Studies
In addition to the study described in Section˜4.2 and Section˜L.4, we ran a supplementary user study to evaluate how explanations and qualities affect users’ trust in model predictions. Each of the 10 questions was presented in three successive stages; after each stage, participants indicated whether they believed the model’s answer was correct or were unsure:
@@
Participants were paid a \$2 base fee and could earn up to \$1 in performance‐based bonuses, which were awarded only during Stage 3 (Explanation + Quality; see Section˜L.3.2).
@@
Table˜24 shows that as users progress from seeing only the model’s answer to viewing explanations and then explanations with quality scores, their unsure rate steadily decreases while user accuracy correspondingly increases, with the largest gains observed when descriptive qualities are provided alongside explanations. These findings support our findings that richer, more interpretable quality signals can meaningfully improve users’ trust calibration.
Appendix M Computational Resources Spent & Total Cost
The complete breakdown of all monetary expenditures across the study is given in Table˜25. In brief, the computational expenses include six hours of A100 GPU time for running the vision‐language models, plus OpenAI API calls for generating predicted answers, explanations, and computing qualities. Human evaluation costs cover two Prolific studies—a one‐stage study as shown in Section˜L.4 and a three‐stage follow‐up from Section˜L.5 (including up to \$1.00 in per‐participation bonuses; see Section˜L.3.2).